a child runs at a constant velocity of 0.90 m/s toward an ice cream truck which is parked 125 m away. If the ice cream truck is only parked for 3 minutes, will the child reach the ice cream truck in time? 


1
Expert's answer
2021-01-10T18:25:36-0500

We can find the time that the child takes to clear the distance to the ice cream truck from the formula:


"t=\dfrac{s}{v}=\dfrac{125\ m}{0.90\ \dfrac{m}{s}}=138.8\ s\cdot(\dfrac{1\ min}{60\ s})=2.3\ min."

As we can see, it takes 2.3 minutes for the child to reach the ice cream truck. Therefore, the child will reach the ice cream truck in time.

Answer:

"t=2.3\ min," the child will reach the ice cream truck in time.